# Little Red Dots and Supermassive Black Hole Seed Formation in Ultralight Dark Matter Halos

### v1

@@ Abstract
We investigate how supermassive black hole (SMBH) seeds form in the early Universe at the centers of ultralight dark matter (ULDM) halos. Focusing on the ULDM Jeans scale, we identify the critical conditions under which high-redshift baryonic gas, strongly confined by central solitonic cores of the halos, undergoes direct and monolithic collapse. The solitonic potential naturally drives rapid inflow and shock heating, allowing the gas to exceed the critical atomic-cooling and fragmentation-suppression threshold of $\sim 3\times 10^{4}{\rm K}$ without invoking an external UV background. We derive semi-analytic relations for the halo mass, soliton mass, baryonic core radius, and thermodynamic state of the gas, including the effects of baryonic contraction. These relations simultaneously determine the minimum and maximum SMBH seed masses as functions of redshift. In this framework, pristine gas clouds that satisfy the temperature threshold collapse without fragmentation, forming SMBH seeds with characteristic masses of order $\sim 10^{5}M_{\odot}$, while systems below the threshold are expected to form compact star clusters instead. Our model also implies an upper limit on the attainable SMBH mass, predicting a maximum mass scale of order $\sim 10^{10}M_{\odot}$, consistent with the most massive quasars observed to date. The ULDM particle mass required to reproduce the inferred seed mass scale, $m\simeq 10^{-22}{\rm eV}$, coincides with the value favored by galactic-scale observations, providing a unified explanation for the characteristic masses of both galactic cores and early SMBH seeds. Our model predicts efficient SMBH seed formation at redshifts $z\gtrsim 10$ and offers a natural interpretation of recently observed little red dots as SMBHs embedded in compact, hot, ionized gas clouds.
@@ I Introduction
The detection of bright quasars less than a billion years after the Big Bang indicates that supermassive black holes (SMBHs) with masses up to $\sim 10^{9}M_{\odot}$ already existed at the redshifts $z\gtrsim 6$, posing a serious challenge to models of black hole seeding and early growth Volonteri (2010); Inayoshi et al. (2020). In stellar-remnant seed scenarios, maintaining nearly continuous and efficient accretion is difficult in the shallow potential wells at high redshift, which motivates massive-seed scenarios where black holes are born with initial masses of $\sim 10^{4}$$10^{6}M_{\odot}$  Volonteri (2010); Latif and Ferrara (2016); Inayoshi et al. (2020).
@@
The direct-collapse black hole (DCBH) scenario assumes that a primordial, metal-free gas cloud undergoes an almost monolithic collapse to form a supermassive star or a quasi-star, which subsequently gives rise to a massive black hole seed Bromm and Loeb (2003); Begelman et al. (2006). The key requirement for this scenario is the suppression of fragmentation, which could be achieved by maintaining the gas temperature at $\sim 10^{4}\,{\rm K}$ and preventing efficient molecular hydrogen cooling, often via a strong Lyman-Werner UV background or dynamical heating Regan and Haehnelt (2009); Shang et al. (2010); Visbal et al. (2014); Inayoshi et al. (2020). Recent observations of little red dots (LRDs) Rusakov et al. (2026) indicate that black hole seeds with masses $M_{bh}\gtrsim 10^{5}M_{\odot}$ formed at the centers of compact ionized hot pristine gas clouds in the early Universe. It is noteworthy that the observed minimum mass of these seeds is comparable to the characteristic minimum mass of galactic cores $\sim 10^{6}M_{\odot}$ Strigari et al. (2008). It is therefore tempting to speculate that these two mass scales are physically connected.
@@
One possible explanation for the characteristic mass scale of galactic cores Lee (2016) is ultralight dark matter (ULDM) Baldeschi et al. (1983); Sin (1994); Lee and Koh (1996); Matos and Guzman (2000); Hu et al. (2000); Boehmer and Harko (2007); Chavanis (2012); Hui et al. (2017a); Matos et al. (2024); Ferreira (2021); Marsh (2016), which has been extensively studied as an alternative to cold dark matter (CDM) and may alleviate small-scale issues of CDM such as the cusp problem. In this model, dark matter consists of extremely light scalar particles with mass $m\simeq 10^{-22}eV$ whose macroscopic de Broglie wavelength leads to quantum pressure supported cores (solitons) embedded in larger halos Schive et al. (2014a); Hui et al. (2017b); Ferreira (2021). These solitonic cores typically have masses $M_{sol}\gtrsim 10^{6}M_{\odot}$ and provide deep, centrally concentrated gravitational potentials that can significantly boost baryonic collapse at high redshift Chiu et al. (2025a), while also helping to alleviate the final parsec problem Koo et al. (2024).
@@
In this paper, we explore a ULDM-assisted SMBH seed formation in which the solitonic core acts as a deep potential well for direct collapse of baryon clouds. The key idea is that baryons gravitationally confined within the soliton can undergo rapid infall accompanied by strong shock heating, thereby driving the gas into a regime where molecular cooling becomes dynamically subdominant, even without invoking an external UV background. We develop semi-analytic scalings for (i) the characteristic halo and soliton properties as a function of ULDM mass $m$ and redshift $z$, and (ii) the thermodynamic state of the collapsing baryons, including the conditions under which compressional heating and shocks overwhelm H2 cooling. We then relate these conditions to the fragmentation criterion to determine the regime in which a global collapse is expected Omukai (2001); Bromm and Loeb (2003); Latif and Ferrara (2016); Inayoshi et al. (2020). In this way, the ULDM soliton provides a natural, purely gravitational route to sustained high inflow rates and elevated temperatures, enabling the formation of supermassive stars with characteristic masses $\sim 10^{5}M_{\odot}$ that can rapidly evolve into SMBH seeds consistent with high-redshift quasar and LRD observations Begelman et al. (2006); Volonteri (2010); Inayoshi et al. (2020).
@@
Section II summarizes the ULDM quantum Jeans scales for halos and the soliton-halo relations used in this work. Section III derives the properties of baryonic core structure inside ULDM solitons, formulates the shock-heating and cooling-imbalance criteria, and presents the fragmentation suppression conditions together with the implied characteristic seed mass scale. Section IV discusses the broader implications of our model for early structure formation, including its physical interpretation of LRDs, predictions for SMBH seed formation across redshift, and the observational consequences and limitations of the scenario.
@@ II Jeans length and halo mass function of Ultralight dark matter
In this section, we present some scales that determine early halos and the corresponding central solitons. The quantum Jeans length scale is given by
@@
where $m_{22}\equiv m/10^{-22}eV$, $h$ is the reduced Hubble constant, and $\Omega_{\rm dm}$ is the dark matter density parameter. The dark matter density at the redshift $z$ is
@@
which implies that ULDM halos can provide sufficiently deep potential wells to confine baryons at high redshift. The dark matter halo mass must satisfy $M_{h}\geq M_{J}$, since the collapsing radius is constrained by $\lambda\geq\lambda_{J}$. The collapse is allowed only in the mass range $M_{\rm upper}\geq M_{h}\geq M_{J}$. The upper bound $M_{\rm upper}$ reflects the observed number density of galactic halos.
@@
From the halo mass, to obtain $M_{sol}$, one may use the empirical formula for the core-halo relation for ULDM Chiu et al. (2025b)
@@
where $\xi^{\prime}(z)=\sqrt{1+z}\,\xi(z)^{1/6}$ and $\xi(z)=(18\pi^{2}+82(\Omega_{m}(z)-1)-39(\Omega_{m}(z)-1)^{2})/\Omega_{m}(z)\simeq 177$ for a large $z$. When $z>7$ (or even $z>1$), the relative change in $\xi(z)$ can be ignored. Taking the halo mass as given above, one has the core soliton mass
@@
with $f_{0}=3.9251$ Hui et al. (2017b). We insert here an extra factor $2\beta/3$ to account for the reduction of the half-mass radius due to the presence of the baryonic matter. This factor $\beta$ may be estimated to be around unity and below we shall set it to be unity for simplicity. To reproduce the observed galactic core sizes, we require $r_{1/2}\simeq{\rm kpc}$, which corresponds to $m\simeq 10^{-22}{\rm eV}$ Schive et al. (2014b). Interestingly, the average dark matter density at the half mass radius
@@
where $\nu\equiv\frac{\delta_{c}}{\sigma}$ with the critical overdensity $\delta_{c}$, and $dn$ is the abundance of halos within a mass interval $dM$. For the function $f(\nu)$ we use the model of Sheth & Tormen Sheth et al. (2001) for simplicity:
@@
To determine $M_{\rm upper}$, one should integrate the HMF and compare the resulting abundance with observations. At a given redshift $z$, we define the maximum halo mass $M_{upper}(z)$ as the mass above which the expected number of halos within the observable Universe over the full sky is exactly unity. To determine this quantity, following Ref. Harrison and Coles (2012), we numerically compute the cumulative comoving number density $n(>M,z)$ from the Sheth–Tormen halo mass function. We then multiply this by the comoving volume element corresponding to a thin redshift slice $dz=0.5$ over the full sky, to obtain the expected number of halos $N(>M,z)$. The maximum halo mass is defined by solving the condition $N(>M_{upper}(z),z)=1$. This construction represents the mass of the most massive halo that is most likely to be observed near that redshift. For simplicity, we use the CDM halo mass function rather than the ULDM one to compute $M_{upper}$, which is a good approximation in the large $M_{h}$ regime. We also assume that the minimum halo mass at $z$ is set by the Jeans mass $M_{J}(z)$. With these assumptions, the minimum and maximum halo masses at a given $z$ can be estimated. Then, from $M_{J}$ and $M_{\rm upper}$, one can obtain the ranges of $\alpha$ and $M_{sol}$ using Eq. (II.7). Fig. 2 shows $M_{J}$ and $M_{upper}$ as functions of $z$ for $m_{22}=(0.1,1,10)$, respectively. As $z$ increases, $M_{upper}$ decreases and the allowed range of $M_{h}$ shrinks.
@@
where $M_{\mathrm{Pl}}$ is the Planck mass. A soliton with a mass exceeding this scale is unstable to gravitational cooling and tends to lose mass  Liebling and Palenzuela (2023), unless it is compressed under extreme conditions such as high-speed collisions or in the very high-redshift Universe. Consequently, we expect the maximum dark matter core mass to be effectively bounded at approximately $M_{max}$. This, in turn, may provide an upper bound on the masses of gas clouds within the core and on the ultimate masses of black hole formed and grown in the gas clouds. Note that $M_{max}$ is usually different from the soliton mass in halos with mass $M_{upper}$.
Figure 2: The minimum mass $M_{J}$ (the thick lines) and the maximum mass $M_{upper}$ (the dashed line) of ULDM halos as functions of the redshift $z$ for $m_{22}=(0.1,1,10)$, respectively. The halo mass decides the core masses of ULDM and those of gas clouds using the core-halo relation. For $M_{upper}$, we used the CDM approximation. At a given $z$, only halos with masses between $M_{J}$ and $M_{upper}$ can exist.
@@ III Baryonic Collapse and Seed-Mass Scale inside ULDM Solitons
Let us now place baryons inside the dark matter soliton. Realistically collapsing atomic gas generically develops a finite-density isothermal core owing to shocks and efficient radiative cooling. Therefore, a pseudo-isothermal profile
@@
is the physically relevant description for the gas cloud  Trinca et al. (2026), where $\rho_{0}$ is the central density and $r_{c}$ is the core radius of the gas cloud. For the peudo-isothermal profile adopted in this work, the baryon mass up to radius $r$ is given by
@@
where $G(x)=x-\arctan(x)$. We adopt this mass profile for initial gas distribution within the ULDM cores. The concentration for the pseudo-isothermal profile $c_{r}\equiv r_{vir}/r_{c}$ can be about $30$ Shapiro et al. (1999), where $r_{vir}$ is the virial radius of the gas clouds. Since we expect baryons to collapse into the gravitational potential well of ULDM cores, for simplicity, we assume $r_{vir}$ is equal to the half-mass radius $r_{1/2}$ of the ULDM soliton cores. Then, $M_{b}(r_{vir})\simeq 0.2M_{sol}/2\simeq 4\pi\rho_{0}r_{c}^{3}G(r_{vir}/r_{c})$, where $0.2$ represents the approximate mass ratio of baryon to dark matter.
From Eq. (II.9), the core radius of baryon becomes
@@
$M_{b}(r_{c})\simeq 4.764\times 10^{4}\alpha\left(\frac{1+z}{m^{2}_{22}}\right)^{3/4}{M_{\odot}}.$ (III.5)
The average baryon density at the radius $r_{c}$ becomes
$\rho_{b}(r_{c})=\frac{3M_{b}(r_{c})}{4\pi r_{c}^{3}}\simeq 6.929\times 10^{-3}\alpha^{4}(1+z)^{3}\,{\rm M}_{\odot}/{\rm pc}^{3}.$ (III.6)
We shall take this gas cloud with $r<r_{c}$ as a baryonic core that triggers collapse via the Jeans instability. The ULDM density at $r=r_{c}$ is much lower than the baryon density and its mass and gravitational contribution may be ignored once the gas core is formed.
The Jeans scale of ULDM determines the minimum halo mass able to collapse at high redshift, while the empirical soliton–halo relation fixes the soliton mass and dark matter core radius. These, in turn, set the baryonic core density, inflow velocity, shock temperature, and accretion rate.
The seed black hole mass may be expressed as
@@
where we adopt a value $f_{bh}\simeq O(0.1)$, motivated by the fact that $M_{bh}$ can grow to as much as about half of the quasi-star mass Hassan et al. (2025); Santarelli et al. (2025). Therefore, the mass of the black hole seeds formed at $z$ satisfies
@@
where $\alpha_{max}$ can be inferred from $M_{upper}$. However, we will show that not all seeds within this mass range can form if the temperature of the gas is not sufficiently high. (See Eq. III.18 below.)
One can also estimate the maximum mass of SMBHs grown from their seeds. If we assume that a seed forms and grows within the most massive physically allowed soliton (i.e., $M_{\rm sol}\simeq M_{\rm max}$) and consumes remaining baryonic mass enclosed within the core, then its mass can be as large as $0.2f_{bh}M_{\rm max}$. This sets the characteristic mass scale of the most massive SMBHs (not just seeds),
@@
Interestingly, from the observational data $M_{bh}^{\max}\simeq 10^{10}M_{\odot}$ Inayoshi et al. (2020), one can again get $m\simeq 10^{-22}eV$ for $f_{bh}=0.1$. Because such massive solitons and the corresponding SMBHs are rare, mergers among them are expected to be infrequent. Consequently, this mass scale is not expected to change significantly due to mergers in cosmic history (see Fig. 3).

### v2

@@ Abstract
We investigate how supermassive black hole (SMBH) seeds form in the early Universe at the centers of ultralight dark matter (ULDM) halos. Focusing on the ULDM Jeans scale, we identify the critical conditions under which high-redshift baryonic gas, strongly confined by central solitonic cores of the halos, undergoes direct and monolithic collapse. The solitonic potential naturally drives rapid inflow and shock heating, allowing the gas temperature to exceed the critical atomic-cooling threshold of $\sim 10^{4}\,{\rm K}$ required for fragmentation suppression without invoking an external UV background. We derive semi-analytic relations for the halo mass, soliton mass, baryonic core radius, and thermodynamic state of the gas, including the effects of baryonic contraction. These relations simultaneously determine the minimum and maximum SMBH seed masses as functions of redshift. In this framework, pristine gas clouds that satisfy the temperature threshold collapse without fragmentation, forming SMBH seeds with characteristic masses of order $10^{5}M_{\odot}$, while systems below the threshold are expected to form compact star clusters instead. Our model also implies an upper limit on the attainable SMBH mass, predicting a maximum mass scale of order $10^{10}M_{\odot}$, consistent with the most massive quasars observed to date. The ULDM particle mass required to reproduce the inferred seed mass scale, $m\simeq 10^{-22}{\rm eV}$, coincides with the value favored by galactic-scale observations, providing a unified explanation for the characteristic masses of both galactic cores and early SMBH seeds. Our model predicts efficient SMBH seed formation at redshifts $z\gtrsim 10$ and offers a natural interpretation of recently observed little red dots as SMBHs embedded in compact, hot, ionized gas clouds.
@@ I Introduction
The detection of bright quasars less than a billion years after the Big Bang indicates that supermassive black holes (SMBHs) with masses up to $\sim 10^{9}M_{\odot}$ already existed at the redshifts $z\gtrsim 6$, posing a serious challenge to models of black hole seeding and early growth Volonteri (2010); Inayoshi et al. (2020). In stellar-remnant seed scenarios, maintaining nearly continuous and efficient accretion is difficult in shallow potential wells at high redshift, which motivates massive-seed scenarios where black holes are born with initial masses of $\sim 10^{4}$$10^{6}M_{\odot}$  Volonteri (2010); Latif and Ferrara (2016); Inayoshi et al. (2020).
@@
The direct-collapse black hole (DCBH) scenario assumes that a primordial, metal-free gas cloud undergoes an almost monolithic collapse to form a supermassive star or a quasi-star which subsequently gives rise to a massive black hole seed Bromm and Loeb (2003); Begelman et al. (2006). The key requirement for this scenario is the suppression of fragmentation, which could be achieved by maintaining the gas temperature at $\sim 10^{4}\,{\rm K}$ and preventing efficient molecular hydrogen cooling, often via a strong Lyman-Werner UV background or dynamical heating Regan and Haehnelt (2009); Shang et al. (2010); Visbal et al. (2014); Inayoshi et al. (2020). Recent observations of little red dots (LRDs) Rusakov et al. (2026) indicate that black hole seeds with masses $M_{bh}\gtrsim 10^{5}M_{\odot}$ formed at the centers of compact, hot, ionized, pristine gas clouds in the early Universe. It is noteworthy that the observed minimum mass of these seeds is comparable to the characteristic minimum mass of galactic cores $\sim 10^{6}M_{\odot}$ Strigari et al. (2008). It is therefore tempting to speculate that these two mass scales are physically connected.
@@
One possible explanation for the characteristic mass scale of galactic cores Lee (2016) is ultralight dark matter (ULDM) Baldeschi et al. (1983); Sin (1994); Lee and Koh (1996); Matos and Guzman (2000); Hu et al. (2000); Boehmer and Harko (2007); Chavanis (2012); Hui et al. (2017a); Matos et al. (2024); Ferreira (2021); Marsh (2016), which has been extensively studied as an alternative to cold dark matter (CDM) and may alleviate small-scale issues of CDM such as the cusp problem. In this model, dark matter consists of extremely light scalar particles with mass $m\simeq 10^{-22}eV$ whose macroscopic de Broglie wavelength leads to quantum pressure supported cores (solitons) embedded in larger halos Schive et al. (2014a); Hui et al. (2017b); Ferreira (2021). These solitonic cores typically have masses $M_{sol}\gtrsim 10^{6}M_{\odot}$ and provide deep, centrally concentrated gravitational potentials that can significantly boost baryonic collapse at high redshift Chiu et al. (2025a), while also helping to alleviate the final parsec problem Koo et al. (2024); Bromley et al. (2024).
@@
In this paper, we explore ULDM-assisted SMBH seed formation in which the solitonic core acts as a deep potential well for direct collapse of baryon clouds. The key idea is that baryons gravitationally confined within the soliton can undergo rapid infall accompanied by strong shock heating, thereby driving the gas into a regime where molecular cooling becomes dynamically subdominant, even without invoking an external UV background. We develop semi-analytic scalings for (i) the characteristic halo and soliton properties as functions of the ULDM mass $m$ and redshift $z$, and (ii) the thermodynamic state of baryons collapsing inside the soliton potential. By applying the zone-of-no-return criterion for the suppression of $H_{2}$ cooling Inayoshi and Omukai (2012), we identify the parameter regime in which fragmentation is avoided and global collapse can occur Omukai (2001); Bromm and Loeb (2003); Latif and Ferrara (2016); Inayoshi et al. (2020). In this way, the ULDM soliton provides a natural, purely gravitational route to sustained high inflow rates and elevated temperatures, enabling the formation of supermassive stars with characteristic masses $\sim 10^{5}M_{\odot}$ that can rapidly evolve into SMBH seeds consistent with high-redshift quasar and LRD observations Begelman et al. (2006); Volonteri (2010); Inayoshi et al. (2020); Bhattacharya et al. (2025).
@@
Section II summarizes the ULDM quantum Jeans scales for halos and the soliton-halo relations used in this work. Section III derives the baryonic core properties inside ULDM solitons and applies the zone-of-no-return criterion for $H_{2}$ cooling suppression to determine the fragmentation suppression conditions and the resulting seed mass scale. Section IV discusses the broader implications of our model for early structure formation, including its physical interpretation of LRDs, the observational consequences and limitations of the scenario.
@@ II Jeans length and halo mass function of Ultralight dark matter
In this section, we present the characteristic scales that determine early halos and the corresponding central solitons. The quantum Jeans length scale is given by
@@
where $m_{22}\equiv m/10^{-22}eV$, $h$ is the reduced Hubble constant, and $\Omega_{\rm dm}$ is the dark matter density parameter. The dark matter density at redshift $z$ is
@@
which implies that ULDM halos can provide sufficiently deep potential wells to confine baryons at high redshift. The dark matter halo mass must satisfy $M_{h}\geq M_{J}$, since a radius $\lambda$ of a ULDM perturbation is constrained by $\lambda\geq\lambda_{J}$. Halo collapse occurs only in the mass range $M_{\rm upper}\geq M_{h}\geq M_{J}$. The upper bound $M_{\rm upper}$ reflects the observed number density of galactic halos.
@@
To obtain $M_{sol}$ from the halo mass one may use the empirical formula for the core-halo relation for ULDM Chiu et al. (2025b)
@@
where $\xi^{\prime}(z)=\sqrt{1+z}\,\xi(z)^{1/6}$ and $\xi(z)=(18\pi^{2}+82(\Omega_{m}(z)-1)-39(\Omega_{m}(z)-1)^{2})/\Omega_{m}(z)\simeq 177$ for a large $z$. When $z>7$ (or even $z>1$), the relative change in $\xi(z)$ can be ignored. Using the halo mass as given above, the core soliton mass becomes
@@
with $f_{0}=3.9251$ Hui et al. (2017b). We insert here an extra factor $2\beta/3$ to account for the reduction of the half-mass radius due to the presence of the baryonic matter. This factor $\beta$ is expected to be of order unity, and we set $\beta=1$ below for simplicity. To reproduce the observed galactic core sizes, we require $r_{1/2}\simeq{\rm kpc}$, which corresponds to $m\simeq 10^{-22}{\rm eV}$ Schive et al. (2014b). Interestingly, the average dark matter density at the half mass radius
@@
where $\nu\equiv\frac{\delta_{c}}{\sigma}$ with the critical overdensity $\delta_{c}$, and $dn$ is the abundance of halos within a mass interval $dM$. For CDM, we adopt the Sheth-Tormen form for $f(\nu)$ Sheth et al. (2001) for simplicity:
@@
To determine $M_{\rm upper}$, one should integrate the HMF and compare the resulting abundance with observations. At a given redshift $z$, we define the maximum halo mass $M_{upper}(z)$ as the mass above which the expected number of halos within the observable Universe over the full sky is exactly unity. To determine this quantity, following Ref. Harrison and Coles (2012), we numerically compute the cumulative comoving number density $n(>M,z)$ from the Sheth–Tormen halo mass function. We then multiply this by the comoving volume element corresponding to a thin redshift slice $dz=0.5$ over the full sky, to obtain the expected number of halos $N(>M,z)$. The maximum halo mass $M_{upper}$ is defined by solving the condition $N(>M_{upper}(z),z)=1$. This construction represents the mass of the most massive halo that is most likely to be observed near that redshift. For simplicity, we use the CDM halo mass function rather than the ULDM one to compute $M_{upper}$, which is a good approximation in the large $M_{h}$ regime. We also assume that the minimum halo mass at $z$ is set by the Jeans mass $M_{J}(z)$. With these assumptions, the minimum and maximum ULDM halo masses at a given $z$ can be estimated. Then, from $M_{J}$ and $M_{\rm upper}$, one can obtain the ranges of $\alpha$ and $M_{sol}$ using Eq. (II.7). Fig. 2 shows $M_{J}$ and $M_{upper}$ as functions of $z$ for $m_{22}=(0.1,1,10)$, respectively. As $z$ increases, $M_{upper}$ decreases and the allowed range of $M_{h}$ shrinks.
@@
where $M_{\mathrm{Pl}}$ is the Planck mass. A soliton with a mass exceeding this scale is unstable to gravitational cooling and tends to lose mass  Liebling and Palenzuela (2023), unless it is compressed under extreme conditions such as high-speed collisions or in the very high-redshift Universe. Consequently, we expect the maximum dark matter core mass to be effectively bounded at approximately $M_{max}$. This, in turn, may provide an upper bound on the masses of gas clouds within the core and on the ultimate masses of black holes formed and grown in the gas clouds. Note that $M_{max}$ is usually different from the soliton mass in halos with mass $M_{upper}$ which is obtained from HMF.
Figure 2: The minimum mass $M_{J}$ (the thick lines) and the maximum mass $M_{upper}$ (the dashed line) of ULDM halos as functions of the redshift $z$ for $m_{22}=(0.1,1,10)$, respectively. The halo mass decides the core masses of ULDM and those of gas clouds using the core-halo relation. For $M_{upper}$, we used the CDM approximation. At a given $z$, only halos with masses between $M_{J}$ and $M_{upper}$ can exist. Smaller $m_{22}$ leads to later halo formation.
@@ III Baryonic Collapse and Seed-Mass Scale inside ULDM Solitons
We now consider baryons embedded within the dark matter soliton. In realistic collapse scenarios, atomic gas generically develops a finite-density isothermal core owing to shocks and efficient radiative cooling Patrick et al. (2023). Therefore, a pseudo-isothermal profile
@@
is the physically relevant description for the gas cloud  Trinca et al. (2026), where $\rho_{0}$ is the central density and $r_{c}$ is the core radius of the gas cloud. For the pseudo-isothermal profile adopted in this work, the baryon mass up to radius $r$ is given by
@@
where $G(x)\equiv x-\arctan(x)$. We adopt this mass profile for initial gas distribution within the ULDM cores. The concentration parameter for the pseudo-isothermal profile $c_{r}\equiv r_{vir}/r_{c}$ is typically about $30$ Shapiro et al. (1999), where $r_{vir}$ is the virial radius of the gas clouds. Since we expect baryons to collapse into the gravitational potential well of ULDM soliton cores, for simplicity, we assume $r_{vir}$ is equal to the half-mass radius $r_{1/2}$ of the ULDM soliton cores. Then, $M_{b}(r_{vir})\simeq 0.2M_{sol}/2\simeq 4\pi\rho_{0}r_{c}^{3}G(r_{vir}/r_{c})$, where $0.2$ represents the approximate mass ratio of baryon to dark matter.
From Eq. (II.9), the baryonic core radius becomes
@@
$M_{b}(r_{c})\simeq 4.764\times 10^{4}\alpha\left(\frac{1+z}{m^{2}_{22}}\right)^{3/4}{M_{\odot}}$ (III.5)
using Eq. (II.8). The average baryon density at the radius $r_{c}$ becomes
$\rho_{b}(r_{c})\simeq\frac{3M_{b}(r_{c})}{4\pi r_{c}^{3}}\simeq 6.929\times 10^{-3}\alpha^{4}(1+z)^{3}\,{\rm M}_{\odot}/{\rm pc}^{3}.$ (III.6)
We shall take this gas cloud with $r\leq r_{c}$ as the initial baryonic core that undergoes subsequent collapse via the Jeans instability. Therefore, $r_{c}$ represents the initial size of the baryonic core that undergoes direct collapse. As the collapse proceeds, the core contracts to smaller radii and its baryon density increases beyond the value given in the above equation. The ULDM density at $r=r_{c}$ is much lower than the baryon density and ULDM mass and its gravitational contribution may be ignored once the gas core is formed.
The Jeans scale of ULDM determines the minimum halo mass able to collapse at high redshift, while the empirical soliton-halo relation fixes the soliton mass and dark matter core radius. These, in turn, set the baryonic core density, inflow velocity, shock temperature, and accretion rate.
The mass of the seed black hole may be expressed as
@@
where we adopt a value $f_{bh}\simeq O(0.1)$, motivated by the fact that $M_{bh}$ can grow to as much as about half of the quasi-star mass Hassan et al. (2025); Santarelli et al. (2025). Therefore, the mass of the black hole seeds formed at $z$ may satisfy
@@
where $\alpha_{max}$ can be inferred from $M_{upper}$. However, we will show that not all seeds within this mass range can form if the temperature of the gas is not sufficiently high. (See Eq. (III.22) below.)
One can also estimate the maximum mass of SMBHs grown from their seeds. If we assume that a seed forms and grows within the most massive physically allowed soliton (i.e., $M_{\rm sol}\simeq M_{\rm max}$) and consumes all remaining baryonic mass enclosed within the core, then its mass can be as large as $0.2f_{bh}M_{\rm max}$. This sets the characteristic mass scale of the most massive SMBHs (not just seeds),
@@
Interestingly, from the observational data $M_{bh}^{\max}\simeq 10^{10}M_{\odot}$ Inayoshi et al. (2020), one again obtains $m\simeq 10^{-22}eV$ for $f_{bh}=0.1$. Because such massive solitons and the corresponding SMBHs are rare, mergers among them are expected to be infrequent. Consequently, this mass scale is not expected to change significantly due to mergers in cosmic history (see Fig. 3).